Parth (parthkohli):

How many four digit numbers are there such that the digits add up to 10 assuming that all digits are allowed to be zero?

OpenStudy (amistre64):

less than 9999 but more than 1000 :)

OpenStudy (amistre64):

1117 1171 1711 7111 1126 1162 1261 1621 2611 1216 1612 2161 6121 6211 2116 6112 etc

Parth (parthkohli):

A more mathematical approach, probably?

mathslover (mathslover):

1000A + 100B + 10C + D = 10 100A + 10B + C + D/10 = 1

mathslover (mathslover):

if the number is ABCD ^

OpenStudy (amistre64):

pfft, that is the most basic mathematical appraoach i know :)

Parth (parthkohli):

I liked what @mathslover said, but that doesn't resolve it. Sorry for taking the medal back. :-(

mathslover (mathslover):

1000A + 100B + 10C + D = 10 ---- (1) A + B/10 + C/100 + D/1000 = 1/100 ---- (2)

Parth (parthkohli):

But we must have four equations, no?

mathslover (mathslover):

No problem :) Parth , thanks though..

mathslover (mathslover):

I am going to subtract the above 2 eqns

OpenStudy (amistre64):

its most likely a counting application with permutations since order matters

OpenStudy (amistre64):

how many basic ways are there to define 4 digits that add up to ten, and then how many ways can each group be sorted thru

Parth (parthkohli):

Can we use combinatorics for this?

mathslover (mathslover):

hmn yes ^^ wait

OpenStudy (amistre64):

combinatorics brings to mind the nCr stuff; i think the nPr stuff might be more applicable

OpenStudy (anonymous):

sorry to correct what @mathslover is doing is a bit wrong /////// A+B+C+D=10 is the requirement///////////

OpenStudy (amistre64):

1117; how many ways are there to permute this? 1126 how many ways? 1135 1144

OpenStudy (amistre64):

we have to account for 0 i spose, which is added computations

mathslover (mathslover):

OH yes @harsh314 thanks a lot

OpenStudy (amistre64):

4+12+12+12 +things with zeros

OpenStudy (amistre64):

118: stuff like 1180, 1108, 1018, but not 0118 127 136 145 19 28 37 46 55

OpenStudy (amistre64):

so permute the 3 groups and times it by 3

OpenStudy (amistre64):

1900, 1090, 1009 looks to be 3 times the 2 groups as well except for the 55 :) 5500, 5050, 5005 is just 3

Parth (parthkohli):

This is very long. :-|

Parth (parthkohli):

Instead of brute-force, can we try something simpler?

mathslover (mathslover):

I got it

OpenStudy (amistre64):

this isnt brute force anymore, its refining

mathslover (mathslover):

Lemme check my method again

mathslover (mathslover):

Forget it . . . I think I must try it again :(

OpenStudy (amistre64):

im to focused on my exam in 30 minutes to think thru this clearerer

mathslover (mathslover):

Parth, do you have answer of it ? ?

Parth (parthkohli):

Nope.

OpenStudy (anonymous):

maybe closer the answer

mathslover (mathslover):

84 numbers ... I think

Parth (parthkohli):

How?

mathslover (mathslover):

I will put up soln here soon... just thinking of more no.

mathslover (mathslover):

can the digits repeat Parth?

Parth (parthkohli):

Yes.

Parth (parthkohli):

Well, a number still has four digits if the digits repeat. :-)

mathslover (mathslover):

I found this on net : http://answers.yahoo.com/question/index?qid=20080330145752AA36Ac2 Now, I am thinking can we use this method to find the answerr?

Parth (parthkohli):

We can think of the ones with zeroes like this: three digit numbers that add up to 10 * 4! two digit numbers that add up to 10 * 4!/2 there can't be 3 or 4 zeroes.

OpenStudy (anonymous):

5193 is the answer or am i wrong as usual

Parth (parthkohli):

Hey wait, did I mention that the first digit can be zero?

OpenStudy (amistre64):

1117 is unique, 4 11xy has 12, and there are 3 ways to chose xy, and no zero groups 4(1)+3(12) 118 is unique, there are 3 ways, and 3 zero groups 1xy has 6 ways, and there are 4 ways to choose xy, and 3 zero groups 3(3)+3(6(4)) 55 is unique, theee is ony 1 way, and 3 zero groups xy 2 ways, there are 4 ways to choose xy, and 3 zero groups 3(1) + 3(2(4)) is there a flaw here?

mathslover (mathslover):

Oh! Is it allowed to have 0910 as the number ?

Parth (parthkohli):

Yes.

OpenStudy (amistre64):

the first digit cant be zero and it still be a 4-digit number tho

OpenStudy (anonymous):

no but a number with first digit as 0 does not make sense look at my answer plz i think it is right

mathslover (mathslover):

^ yeah after that 0910 is not a four digit number , it will be three digit number ..

OpenStudy (amistre64):

have fun, i gotta run ;)

Parth (parthkohli):

I know... the question is making such an assumption, so...

mathslover (mathslover):

@amistre64 ;)

mathslover (mathslover):

well there can't be more than 186 numbers.. I think.

mathslover (mathslover):

^ 196

mathslover (mathslover):

oh sorry 186

Parth (parthkohli):

Is it 186?

OpenStudy (anonymous):

more than 577

mathslover (mathslover):

yes it is 186 @ParthKohli

Parth (parthkohli):

The answer is 186?

OpenStudy (anonymous):

(0..9).each do |a| (0..9).each do |b| (0..9).each do |c| (0..9).each do |d| count = count + 1 if (a + b + c + d) == 10 end end end end puts count gave me 282 as answer, I don't know a mathematical approach though :)

mathslover (mathslover):

yes @ParthKohli

Parth (parthkohli):

186 when 0 is allowed or not?

mathslover (mathslover):

allowed

Parth (parthkohli):

How? :S

Parth (parthkohli):

282 and 186 are both not right :-|

OpenStudy (anonymous):

i think i got the answer.............

OpenStudy (anonymous):

282 was with leading zeros, without it would be 219

OpenStudy (anonymous):

408 without the zeros

OpenStudy (anonymous):

468 with zeroes

OpenStudy (anonymous):

plz do reply if wrong.......

OpenStudy (anonymous):

@ParthKohli

Parth (parthkohli):

I'd check

Parth (parthkohli):

Done! 13 C 3 = 286

OpenStudy (anonymous):

yes i got this too a silly error how did you do it............

OpenStudy (anonymous):

my method is tooo long.......

OpenStudy (anonymous):

@ParthKohli

Parth (parthkohli):

... 286 is the answer

OpenStudy (anonymous):

i too got the same but by long method how did you get it .........

OpenStudy (anonymous):

I still find it odd that my bruteforce method which calculates the sum of the digits from 0 upto 9999 is off by 4

Parth (parthkohli):

We wanted the number of non negative numbers \(a_1,a_2,a_3,a_4\) such that \(a_1 + a_2 + a_3 + a_4 = 10\). A formula says that it would be \(\binom{10 + 4 - 1}{4 - 1}\)

OpenStudy (anonymous):

oh so no logic behind it as such..............

OpenStudy (anonymous):

But bruteforcing every digit from 0 to 9999 gives 282 though, shouldn't they be the same?

Parth (parthkohli):

I don't know. :-|

OpenStudy (anonymous):

Weird :l

OpenStudy (anonymous):

my method was long and a bit calculation mistake.............. case1:::'1" as the first digit ,,,,,, the following subcases arise (a) .....with 2 zeroes ........9 has to be the sum of the remaining three numbers so total possible combinations are 3. (b) .......with 1 zero we have 810,720,630,540......all these have 6 combinations each so total 24 (c) .......with no zeroes.....711(3combinations),621(6),531(6),522(3),432(6),441(3),333(1) so total of 28 combinations hence for the 1st case we have 45 combinations similarly we can have others tooo..........and then add all........we get 286

OpenStudy (anonymous):

but it would be great if any einstein of the present century could explain how @ParthKohli 's method came into being.............

OpenStudy (anonymous):

@satellite73

OpenStudy (amistre64):

works in google chrome :) <html> <body> <input type=button value="AddUp" onClick="qwe()"><br> 0000: <input id="txtOut0"><br> 1000: <input id="txtOut1"> </body> <script language=javascript> var a=0,b=0,c=0,d=0 var count=0 function qwe() {for (var x=1000; x<=10000; x=x+1) {if (a+b+c+d == 10){count=count+1} d=d+1; if (d>10) {d=0; c=c+1} if (c>10){c=0; b=b+1} if (b>10){b=0; a=a+1} if (a >9){x=10001} } document.getElementById("txtOut0").value=count //resets a=1,b=0,c=0,d=0, count=0 for (var x=1000; x<=10000; x=x+1) {if (a+b+c+d == 10){count=count+1} d=d+1; if (d>10) {d=0; c=c+1} if (c>10){c=0; b=b+1} if (b>10){b=0; a=a+1} if (a >9){x=10001} } document.getElementById("txtOut1").value=count } </script> </html>

OpenStudy (amistre64):

i should prolly include an output such that we can verify its a+b+c+d parts to make sure its operating as expected

OpenStudy (anonymous):

I'm working on one :)

OpenStudy (amistre64):

gotta weed out a bug :) 0,0,0,10 0,0,1,9 0,0,2,8 0,0,3,7 0,0,4,6 0,0,5,5 0,0,6,4 0,0,7,3 0,0,8,2 0,0,9,1 0,0,10,0 0,1,0,9 0,1,1,8

OpenStudy (amistre64):

if abcd = 10 then reset to 0, not if greater than 10 .... lol

OpenStudy (amistre64):

im missing 9001, 9010, 9100 from the end of my 0xxx for some reason 279+3 = 282 otherwise its 219 if we use actual 4-digit numbers :)

OpenStudy (anonymous):

282 is correct :)